Circular No. 5601

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992ap IN UGC 10430 A. V. Filippenko, T. Matheson and L. C. Ho, University of California,

Berkeley, report that an uncalibrated CCD spectrum (range 400-700 nm) obtained on Sept. 2 UT with the Shane 3-m reflector at the Lick Observatory confirms that SN 1992ap is of type Ia, ∼ 1-2 months past maximum brightness. However, the declination of UGC 10430 is +41^{o}35′25′′ (equinox 1950.0), rather than the +41^{o}25′25′′ that was stated on IAUC 5573.

SUPERNOVA 1992av IN ANONYMOUS GALAXY S. Benetti, F. Patat and M. Turatto, Asiago Astrophysical Observa-

tory, report that they obtained a low-resolution spectrum of this faint object on Sept. 1.92 UT with the Boller & Chivens spectrograph (+ CCD) on the 1.8-m telescope at Cima Ekar. The object is indeed a supernova with a relatively blue continuum, broad emissions measured at 498, 534, 569 and 638 nm and two absorptions at 617 and 668 nm. There seems to be a strong resemblance to the spectrum of normal type Ia supernovae (e.g., SN 1989B) ∼ 2 weeks past maximum. Assuming an average expansion velocity for the supernova’s envelope, the redshift deduced from the Si II doublet is z ∼ 0.09.

1992 HE

Improved orbital elements from MPC 20644:

Epoch = 1992 June 27.0 TT

T = 1992 June 21.184 TT  ω = 262.^{o}600
e = 0.57177  Ω = 27.321 }2000.0
q = 0.95963 AU  i = 37.370
a = 2.24094 AU n^{o} = 0.293805 P = 3.35 years
1992 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Aug. 26 4^{h}18.^{m}90 − 6^{o}45.^{′}6 0.742 1.269 91.^{o}5 52.^{o}7 15.8
Sept. 5 4 09.06 − 2 09.7 0.707 1.347 102.0 47.1 15.7
15 3 52.74 + 2 45.4 0.674 1.426 114.7 39.8 15.5
25 3 28.83 + 7 56.4 0.653 1.506 129.6 30.9 15.3
Oct. 5 2 57.81 +13 02.2 0.654 1.586 146.1 20.6 15.1
15 2 22.94 +17 26.1 0.687 1.665 162.6 10.3 15.0
25 1 49.45 +20 39.3 0.754 1.743 171.2  5.0 15.0

1992 September 2 Brian G. Marsden